Material, Labor, and Variable Overhead Variances
The following summarized manufacturing data relate to Thomas Corporation's April operations, during which 2,000 finished units of product were produced. Normal monthly capacity is 1,100 direct labor hours.

Standard Units Costs Total Actual Costs
Direct material
Standard (2 lb. @ $18.00/lb.) $36
Actual (4,200 lb. @ $19/lb.) $79,800
Direct labor
Standard (0.5 hr. @ $33/hr.) $16.50
Actual (950 hrs. @ $32.20/hr.) 30,590
Variable overhead
Standard (0.5 hr. @ $15/hr.) $7.50
Actual - 15,450
Total $60 $125,840

Determine the following variances:

Do not use negative signs with any of your answers. Next to each variance answer, select either "F" for Favorable or "U" for Unfavorable.

Materials Variances
Actual cost: Answer
Split cost: Answer
Standard cost: Answer
Materials price Answer Answer
Materials efficiency Answer Answer
Labor Variances
Actual cost: Answer
Split cost: Answer
Standard cost: Answer
Labor rate Answer Answer
Labor efficiency Answer Answer
Variable Overhead Variances
Actual cost: Answer
Split cost: Answer
Standard cost: Answer
Variable overhead spending Answer Answer
Variable overhead efficiency Answer Answer
Answer #1

1 I B C Materials Variance 2 Actual cost $ 79,800 3 Split Cost $ 75,600 4 Standard cost $ 72,000 5 Material price $ 4,200 U 6

Material price variance = (Actual price - Standard price) Actual quantity Material price variance = ($19 - $18)*4,200 Materi

1 Bc Materials Variance 2 Actual cost 79800 3 Split Cost -4200*18 4 Standard cost = 2000*36 5 Material price =+B2-B3 6 Materi